OpenStudy (idealist10):

Use the principal of superposition to find a particular solution of y"+2y'+10y=4+26x+6x^2+10x^3+e^(x/2).

OpenStudy (idealist10):

@freckles @mathstudent55 @phi @Compassionate @shifuyanli @paki @campbell_st

OpenStudy (freckles):

ok so we want to look at solving both of the following for the particulars: \[y''+2y'+10y=e^\frac{x}{2} \\ y''+2y'+10y=4+26x+6x^2+10x^3\]

OpenStudy (freckles):

and then take those particulars and add them together to find the particular of our differential equation above

OpenStudy (idealist10):

So that's the method for using the principal of superposition?

OpenStudy (freckles):

yep if y1 and y2 are solutions to the dq, then y1+y2 is also a solution to the dq

OpenStudy (idealist10):

So what do I do with those ^ above?

OpenStudy (freckles):

find the particular solutions for both

OpenStudy (freckles):

I would probably use y=Aexp(x/2) for the first one and probably use y=Kx^3+Bx^2+Cx+D for the second one as my choice for particular solutions of course you need to find A,K,B,C, and D

OpenStudy (idealist10):

Let me do them real quick, please wait.

OpenStudy (michele_laino):

I think that for second one we have to choose a polynomial of fifth degree, namely: \(P(x)=a_0+a_1x+a_2x^2+a_3x^3+a_4x^4+a_5x^5\)

OpenStudy (freckles):

I still think we only need to do a 3rd degree polynomial since we have that that right hand side of our differential equation is a 3rd degree polynomial http://tutorial.math.lamar.edu/Classes/DE/UndeterminedCoefficients.aspx

OpenStudy (idealist10):

@Michele_Laino , what's your method and why 5th degree?

OpenStudy (freckles):

you can do a 5th degree polynomial but I feel like you would be wasting your time just a bit @Idealist10 since you will get the coefficients of x^4 and x^5 are zero

OpenStudy (michele_laino):

I choose a fifth degree polynomial, since after two derivatives, as indicated by your ODE, I get a polynomial of third degree

OpenStudy (michele_laino):

and the right side is a third degree polynomial

OpenStudy (michele_laino):

If we compute the second derivative we get: \[y'' = 2{a_2} + 6{a_3}x + 12{a_4}{x^2} + 20{a_5}{x^3}\]

OpenStudy (freckles):

but you have no x^4 and x^5 's on the right hand side

OpenStudy (freckles):

the coefficients would have to be 0 of x^4 and x^5

OpenStudy (michele_laino):

and if \(a_4,a_5\) are both zero, then we can not have a third degree polynomial

OpenStudy (idealist10):

Okay, so I've got the right answer. But what if we just use \[y _{p}=A+Bx+Cx^2+Dx^3+Ae ^{\frac{ x }{ 2 }}\] as the substitution instead of doing it 2 times like @freckles suggested at the beginning?

OpenStudy (freckles):

we could have but I just wanted to use the principle they mentioned

OpenStudy (jango_in_dtown):

Lets use shortcut methods in evaluating the particular solutions

OpenStudy (freckles):

the principle of superposition

OpenStudy (michele_laino):

The best way is to write the solution with both methods

OpenStudy (freckles):

\[y''+p(x)y'+q(x)y=g(x)\] basically it says if g(x)=g1+g2 then \[y''+py'+qy=g_1 \text{ will give solution } y_1 \\ \text{ and } y''+py'+qy=g_2 \text{ will give solution } y_2 \\ \text{ then the solution to } \\ y''+py'+qy=g \text{ is } y_1+y_2\]

OpenStudy (idealist10):

Now I don't like to use the principal of superposition. It's kinda irritating. :)

OpenStudy (michele_laino):

I got this: \(P(x)=2x+x^3\)

OpenStudy (jango_in_dtown):

check if you can compute the remaining part. or else I will compute

OpenStudy (idealist10):

Thanks a lot for the help, guys!